OpenStudy (anonymous):

cotx/tanx=1-sin^2x/1-cos^2x

OpenStudy (anonymous):

need to prove left side = right side

OpenStudy (anonymous):

what is cot x? what is tan x? do you know their identity?

OpenStudy (anonymous):

its actually theta

OpenStudy (anonymous):

theta???? you mean tan (theta) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it doesn't really matter... it could be tan(elephant) cos(monkey) sin(SQUARE) its the variable, you can assign it to anything if i said sin(ANT) and ANT is 0 then sin(0) =0 do you see?

OpenStudy (anonymous):

so do you know their identities?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, the identities of ... \[\tan(x)=\frac{ \sin(x) }{ \cos(x) }\] \[\cot(x)=\frac{ 1 }{ \tan(x) }\] and \[\cos ^{2}(x)+\sin ^{2}(x)=1\] so using these you should be able to work them out, ask me if you get stuck

OpenStudy (anonymous):

 i dont know what to do after